OpenStudy (anonymous):

Given: costheta= -4/5, sin x =-12/13 ,theta is in the third quadrant, and x is in the fourth quadrant; evaluate cos 2x. -119/169 119/169 -11/13

OpenStudy (anonymous):

you don't need the sine part

OpenStudy (anonymous):

then how do I start

terenzreignz (terenzreignz):

Two variables?

OpenStudy (anonymous):

\[\cos(2x)=2\cos^2(x)-1\]

OpenStudy (anonymous):



OpenStudy (anonymous):

oh i see there is a \(\theta\) and an \(x\)

OpenStudy (anonymous):

costheta= -4/5 I need cosx

OpenStudy (anonymous):

you are asked for \(\cos(2x)\) where \(\sin(x)=-\frac{12}{13}\)

OpenStudy (anonymous):

\[\cos(2x)=1-2\sin^2(x)\]

OpenStudy (anonymous):

When I do it I get 169

OpenStudy (anonymous):

a denominator of 169

OpenStudy (anonymous):

there is no \(\theta \) in your question

OpenStudy (anonymous):

ok, I'll do it for you, give me a second, this is an easy question

OpenStudy (anonymous):

yea I had to put theta for that

OpenStudy (anonymous):

satellite, you don't know what you're doing..

OpenStudy (anonymous):

\[\cos(x)=1-2\sin^2(x)\] \[\cos(x)=1-(\frac{-12}{13})^2\]\

OpenStudy (anonymous):

there is your 169

OpenStudy (anonymous):

\[\cos(2x)=1-(\frac{12}{13})^2\] \[\cos(2x)=1-\frac{144}{169}\]

OpenStudy (anonymous):

\[\cos(2x)=\frac{25}{169}\]

OpenStudy (anonymous):

I know, thats what I get, but it isnt a choice

OpenStudy (anonymous):

your question asks for \(\cos(2x)\) and doesn't mention any \(\theta\)

OpenStudy (anonymous):

I know but that is what is already given

OpenStudy (anonymous):

oh damn i made a mistake!! sorry

OpenStudy (anonymous):

I'm too lazy to do it but you need to draw the triangles and either use theta or x for the angle

OpenStudy (anonymous):

lol its okay, Ive made plenty

OpenStudy (anonymous):

\[\cos(2x)=1-2\sin^2(x)\] \[\cos(2x)=1-2\times (\frac{12}{13})^2\] \[cos(2x)=1-2\times \frac{144}{169}\]

OpenStudy (anonymous):

\[\cos(2x)=1-\frac{288}{169}\]\[\cos(2x)=-\frac{119}{169}\]

OpenStudy (anonymous):

Lol this is a lot of work, can you use this same formula for tan2Theta

OpenStudy (anonymous):

nvm. I found the formula, Thank you!